# Diffuse Neutrino Background from Magnetorotational Stellar Core Collapses

### v1

@@ Abstract
A statistically significant detection of the diffuse supernova neutrino background (DSNB) is around the corner. To this purpose, we assess the contribution to the DSNB of magnetorotational collapses of massive stars, relying on a suite of state-of-the-art three-dimensional neutrino-magnetohydrodynamic simulations. We find that neutrinos from magnetorotational core collapses boost the high-energy tail of the DSNB spectrum, similar to what is expected from neutrino-driven black hole-forming collapses. The latest data from the Super-Kamiokande Collaboration can already exclude that more than $13\%$ of all collapsing massive stars undergo magnetorotational collapses under optimistic assumptions. A DSNB detection at $3\sigma$ could take place up to $4$ yr earlier at Super-Kamiokande-Gadolinium or JUNO if the fraction of magnetorotational collapses should be larger than $10\%$. Fascinatingly, if the fraction of magnetorotational stellar collapses should be larger than $7\%$, Hyper-Kamiokande could measure such a fraction at $3\sigma$ after $20$ yr of DSNB data taking. The combination of DSNB and electromagnetic data has the potential to resolve the degenerate contributions from magnetorotational and neutrino-driven black hole forming collapses, providing crucial insight on the properties of the population of collapsing massive stars.
@@ I Introduction
In this paper, we investigate the contribution to the diffuse neutrino emission from magnetorotational core collapses of massive stars, leading to protomagnetars and spinars.  [^1] Note that we distinguish between magnetorotational core collapses leading to the formation of a protomagnetar or a spinar (if the central compact object is a BH). The term “spinar” was first introduced in Ref. [34] in the context of active galactic nuclei, but it is now commonly adopted to refer to a fast rotating magnetized relativistic object, in quasi-equilibrium due to the balance between the centrifugal and gravitational forces, which undergoes a two-stage collapse, first to a protoneutron star and, subsequently, to form a BH [35, 36, 37, 38]. Magnetorotational core collapses are considered to be the progenitors of superluminous SNe and certain gamma-ray bursts [39, 40]. To this purpose, we model the neutrino emission from the (transiently stable) protomagnetars relying on three-dimensional neutrino-magnetohydrodynamic simulations of the same models studied in 2D in Ref. [40]. Since the rate of protomagnetars and spinars is unknown, we keep it as a free parameter and investigate whether the neutrino emission from magnetorotational core collapses could lead to a DSNB contribution which is degenerate with respect to the uncertainties in the CCSN rate and fraction of neutrino-driven BH-forming collapses.
@@ II.2 Protomagnetar and spinar models
Figure 1 shows the angle-averaged (1D equivalent) neutrino emission properties for the representative protomagnetar model with mass $M=17M_{\odot}$ (model A17; note that the sudden termination of the neutrino signal at about $4$ s corresponds to the end of the simulation time for this model and not to BH formation) and the spinar model with mass 26 $M_{\odot}$ (model A26; the signal terminates about $1.5$ s after core bounce due to BH formation). These models have neutrino luminosity $\text{L}_{\nu}\gtrsim 10^{52}$ erg s-1 for all flavors, with the typical neutrino emission lasting for several seconds. The neutrino mean energies for the different flavors exhibit a characteristic hierarchy: $\langle\varepsilon_{\nu_{e}}\rangle<\langle\varepsilon_{\bar{\nu}_{e}}\rangle<\langle\varepsilon_{\nu_{x}}\rangle$. The large mean energy of the non-electron flavor neutrinos (which exceeds $20$ MeV for a significant part of the emission) is due to the compact protoneutron star, that facilitates the neutrino decoupling at densities higher than in the neutrino-driven CCSN models. The small offset (of a few MeV) in the heavy-neutrino mean energies is produced by the development of steep density and temperature gradients at the typical densities where the transition region from diffusion to free streaming is located. As a result, the flux factors (i.e., the ratio of flux-to-neutrino energy) grow over a narrow radial range. The region where $\nu_{x}$’s decouple from matter and the other neutrino flavors coincides with a sharp drop of the gas temperature, such that even a small radial difference has an impact on the mean energies of the non-electron neutrino flavors. The detailed treatment of the pair processes may have also some (small) impact on the computed heavy-neutrino mean energy.
@@ III.2 Contribution to the diffuse neutrino background from protomagnetars and spinars
Assuming that our protomagnetar model A17 or our spinar model A26 are representative of the population of massive stellar collapses, the top panel of Fig. 2 shows the corresponding $\bar{\nu}_{e}$ fluxes.  [^2] We have verified that the neutrino spectral energy distributions of our suite of magnetorotational collapse models follow the parametrization of Refs. [49, 50]. One can see that the $\bar{\nu}_{e}$ spectra from protomagnetars and spinars exhibit an energy tail with slope comparable to the one of the neutrino-driven BH-forming collapse model, but peaking at $E\simeq 10$ MeV in the source frame.
@@ V.1 Near term detection prospects
The increase in Gadolinium doping in Super-Kamiokande is expected to result in a larger event statistics also for what concerns the DSNB [76, 77]. To illustrate the short-term potential of Super-Kamiokande-Gadolinium, we investigate the exposure that one would need to detect the DSNB and infer the fraction of magnetorotational collapses.
@@
For a fiducial volume of $22.5$ kT, $65\%$ detection efficiency, and the backgrounds reported for the Super-Kamiokande-VI and Super-Kamiokande-VII runs [27], we compute the ratio between the signal events ($S$) and the Poissonian statistical fluctuations of the total number of events, i.e. signal plus background ($B$). For our baseline analysis, the fraction of magnetorotational collapses is the only free parameter for a fixed star formation history (i.e., for a given set of values $\dot{\rho}_{0}$, $\alpha$, $\beta$ and $\gamma$; see Table 1)–whereas the CCSN rate ( Eq. 4) and the fraction of neutrino-driven BH-forming collapses ($f_{\nu\text{BH}}=0.3$) are assumed to be known.
@@
Hyper-K-Gd $374$ $0.67$ 9.49.49.49.431.431.431.431.4 $17.5$ $22.6$ $18.0$ ($10.5$  [^3] This value corresponds to the scenario in which neutral current backgrounds induced by atmospheric neutrinos are negligible.)
@@ V.2 Long term detection prospects
A better assessment on the contribution from protomagnetars and spinars to the DSNB is expected from the upcoming water Cherenkov neutrino telescope Hyper-Kamiokande [73]. We consider an experimental configuration with two tanks of $187$ kT each and detection efficiency of $67\%$, resulting from adding $\sim 0.1\%$ of GdCl3 to water (cf. Table 2 for a summary on the expected number of events). Among the different backgrounds expected in Hyper-Kamiokande, neutral-current atmospheric background events are particularly relevant [80, 81], and the possibility of reducing them to negligible level is under investigation [82].
@@
$\frac{S(f_{\text{MR}})-S(f_{\text{MR}}=0)}{\sqrt{S(f_{\text{MR}}=0)+B}}=3\,.$ (10)
One can see that a fraction of magnetorotational collapses smaller than $7\%$ could be measured at $3\sigma$ after $20$ yr of data taking, assuming that the fraction of neutrino-driven BH-forming collapses is known and the neutral current atmospheric background is efficiently tagged.
Figure 7: Hyper-Kamiokande sensitivity at $3\sigma$ to the fraction of magnetorotational core collapses ($f_{\text{MR}}$) as a function of the exposure. The solid line corresponds to negligible neutral-current backgrounds induced by atmospheric neutrinos. The dashed line considers the fiducial backgrounds from Ref. [83]. The dark shaded band represents the impact of the uncertainties on the CCSN rate and the fraction of neutrino-driven BH-forming collapses. After $20$ yr exposure, Hyper-Kamiokande-Gadolinium will be sensitive to $f_{\rm{MR}}\gtrsim 7\%$.
The upcoming liquid Argon neutrino experiment DUNE [75], being sensitive to the $\nu_{e}$ component, could also add positively to this picture, although with smaller statistics (cf. Table 2). We estimate that a ratio $S/\sqrt{S+B}=3$ (i.e. to reject the background-only hypothesis at $3\sigma$ level) could only be achieved in $10$ yr at this detector, if $f_{\text{MR}}\gtrsim 0.27$–assuming $100\%$ detection efficiency and the backgrounds from Ref. [83]. The difference between the mean energies of electron and non-electron flavor neutrinos is larger than the one for antineutrinos. This could potentially disentangle some of the existing uncertainties in the picture, especially for what concerns the flavor conversion physics and the fraction of neutrino-driven BH-forming collapses [29].
@@ VII Conclusions and outlook
As the detection of the DSNB approaches, it is timely to address whether the diffuse neutrino emission from magnetorotational core collapses, leading to the formation of protomagnetars or spinars, could contribute to the DSNB. We rely on a suite of seven 3D special relativistic neutrino-magnetohydrodynamic simulations with different ZAMS mass, inspired by the 2D equivalent models presented in Ref. [40], to investigate the related neutrino properties and the associated contribution to the DSNB. Our models predict large neutrino luminosities $\gtrsim 10^{52}$ ergs-1 for several seconds and large mean energies for non-electron neutrinos and antineutrinos ($\langle\varepsilon_{\nu_{x}}\rangle\gtrsim 20$ MeV). Such neutrino properties suggest that magnetorotational core collapses could provide a non-negligible contribution to the high-energy DSNB tail (above $5$ MeV).
@@
Given its larger size, the upcoming Hyper-Kamiokande-Gadolinium neutrino telescope could measure a fraction of magnetorotational collapses larger than $7\%$ at $3\sigma$ after $20$ yr of data-taking, if meanwhile we can infer the fraction of (neutrino-driven) BH-forming collapses and pin down the uncertanties on the CCSN rate through electromagnetic observations. This measurement would further benefit from improved background reduction and detection efficiency.

### v2

@@ Abstract
A statistically significant detection of the diffuse supernova neutrino background (DSNB) is around the corner. To this purpose, we assess the contribution to the DSNB of magnetorotational collapses of massive stars, relying on a suite of state-of-the-art three-dimensional neutrino-magnetohydrodynamic simulations. We find that neutrinos from magnetorotational core collapses boost the high-energy tail of the DSNB spectrum, similar to what is expected from neutrino-driven black hole-forming collapses. The latest data from the Super-Kamiokande Collaboration can already exclude that more than $13\%$ of all collapsing massive stars undergo magnetorotational collapses under optimistic assumptions. A DSNB detection at $3\sigma$ could take place up to $4$ yr earlier at Super-Kamiokande-Gadolinium or JUNO if the fraction of magnetorotational collapses should be larger than $10\%$. Fascinatingly, if the fraction of magnetorotational stellar collapses should be larger than $19\%$ ($13\%$), Hyper-Kamiokande could measure such a fraction at $3\sigma$ after ($10$ yr) $20$ yr of DSNB data taking. The combination of DSNB and electromagnetic data has the potential to resolve the degenerate contributions from magnetorotational and neutrino-driven black hole-forming collapses, providing crucial insight on the properties of the population of collapsing massive stars.
@@ I Introduction
In this paper, we investigate the contribution to the diffuse neutrino emission from magnetorotational core collapses of massive stars, i.e. from the collapse of rapidly rotating stars for which the rotation and magnetic fields are relevant. We differentiate between two magnetorotational collapse scenarios according to the nature of their remnant: a protomagnetar or a spinar. We refer to rapidly rotating protoneutron stars with strong magnetic fields as protomagnetars. We adopt the term spinar for a fast rotating magnetized relativistic object, in quasi-equilibrium due to the balance between the centrifugal and gravitational forces, which undergoes a two-stage collapse, first to a protoneutron star and, subsequently, to form a BH [34, 35, 36, 37]  [^1] The term “spinar” was first introduced in Ref. [38] in the context of active galactic nuclei.
Magnetorotational core collapses are considered to be the progenitors of superluminous SNe and certain gamma-ray bursts [39, 40]. To this purpose, we model the neutrino emission from the (transiently stable) protomagnetars relying on three-dimensional neutrino-magnetohydrodynamic simulations of the same models studied in 2D in Ref. [40]. Since the rate of protomagnetars and spinars is unknown, we keep it as a free parameter and investigate whether the neutrino emission from magnetorotational core collapses could lead to a DSNB contribution which is degenerate with respect to the uncertainties in the CCSN rate and fraction of neutrino-driven BH-forming collapses.
@@ II.2 Protomagnetar and spinar models
Figure 1 shows the angle-averaged (1D equivalent) neutrino emission properties for the representative protomagnetar model with mass $M=17M_{\odot}$ (model A17; note that the sudden termination of the neutrino signal at about $4$ s corresponds to the end of the simulation time for this model and not to BH formation) and the spinar model with mass 26 $M_{\odot}$ (model A26; the signal terminates about $1.5$ s after core bounce due to BH formation). These models have neutrino luminosity $\text{L}_{\nu}\gtrsim 10^{52}$ $\cdots$-1 for all flavors, with the typical neutrino emission lasting for several seconds. The neutrino mean energies for the different flavors exhibit a characteristic hierarchy: $\langle\varepsilon_{\nu_{e}}\rangle<\langle\varepsilon_{\bar{\nu}_{e}}\rangle<\langle\varepsilon_{\nu_{x}}\rangle$. The large mean energy of the non-electron flavor neutrinos (which exceeds $20$ MeV for a significant part of the emission) is due to the compact protoneutron star, that facilitates the neutrino decoupling at densities higher than in the neutrino-driven CCSN models. The small offset (of a few MeV) in the heavy-neutrino mean energies is produced by the development of steep density and temperature gradients at the typical densities where the transition region from diffusion to free streaming is located. As a result, the flux factors (i.e., the ratio of flux-to-neutrino energy) grow over a narrow radial range. The region where $\nu_{x}$’s decouple from matter and the other neutrino flavors coincides with a sharp drop of the gas temperature, such that even a small radial difference has an impact on the mean energies of the non-electron neutrino flavors. The detailed treatment of the pair processes may have also some (small) impact on the computed heavy-neutrino mean energy.
@@ III.2 Contribution to the diffuse neutrino background from protomagnetars and spinars
Assuming that our protomagnetar model A17 or our spinar model A26 are representative of the population of massive stellar collapses, the top panel of Fig. 2 shows the corresponding $\bar{\nu}_{e}$ fluxes.  [^2] We have verified that the neutrino spectral energy distributions of our suite of magnetorotational collapse models follow the parametrization of Refs. [49, 50]. One can see that the $\bar{\nu}_{e}$ spectra from protomagnetars and spinars exhibit an energy tail with slope comparable to the one of the neutrino-driven BH-forming collapse model. However, for magnetorotational core collapses, the energy distributions peak at $E\simeq 10$ MeV in the source frame.
@@ V.1 Near term detection prospects
The increase in Gadolinium doping in Super-Kamiokande has resulted in a larger event statistics also for what concerns the DSNB [76, 77]. To illustrate the short-term potential of Super-Kamiokande-Gadolinium, we investigate the exposure that one would need to detect the DSNB and infer the fraction of magnetorotational collapses.
@@
For a fiducial volume of $22.5$ kT, $55\%$ detection efficiency  [^3] The detection efficiency is the result of a $75\%$ neutron-capture efficiency for a $0.03\%$ of Gd-loaded water and a $74\%$ selection efficiency [25]., and the backgrounds reported for the Super-Kamiokande-VI and Super-Kamiokande-VII runs [27], we compute the ratio between the signal events ($S$) and the Poissonian statistical fluctuations of the total number of events, i.e. signal plus background ($B$). For our baseline analysis, the fraction of magnetorotational collapses is the only free parameter for a fixed star formation history (i.e., for a given set of values $\dot{\rho}_{0}$, $\alpha$, $\beta$ and $\gamma$; see Table 1)–whereas the CCSN rate (see Eq. 4) and the fraction of neutrino-driven BH-forming collapses ($f_{\nu\text{BH}}=0.3$) are assumed to be known.
@@
Hyper-K-Gd $187$ $0.55$ 9.49.49.49.431.431.431.431.4 $14.3$ $18.6$ $18.0$ ($10.5$  [^4] This value corresponds to the scenario in which neutral current backgrounds induced by atmospheric neutrinos are negligible.)
@@ V.2 Long term detection prospects
[…]
@@
$\frac{S(f_{\text{MR}})-S(f_{\text{MR}}=0)}{\sqrt{S(f_{\text{MR}})+B}}=3\,.$ (10)
One can see that a fraction of magnetorotational collapses smaller than $13\%$ ($19\%$) could be measured at $3\sigma$ after $20$ yr (10 yr) of data taking, assuming that the fraction of neutrino-driven BH-forming collapses is known and the neutral current atmospheric background is efficiently tagged.
Figure 7: Hyper-Kamiokande sensitivity at $3\sigma$ to the fraction of magnetorotational core collapses ($f_{\text{MR}}$) as a function of the exposure. The solid line corresponds to negligible neutral-current backgrounds induced by atmospheric neutrinos. The dashed line considers the fiducial backgrounds from Ref. [83]. The dark shaded band represents the impact of the uncertainties on the CCSN rate and the fraction of neutrino-driven BH-forming collapses. After $20$ yr exposure, Hyper-Kamiokande-Gadolinium will be sensitive to $f_{\rm{MR}}\gtrsim 13\%$
The upcoming liquid Argon neutrino experiment DUNE [75], being sensitive to the $\nu_{e}$ component, could also add positively to this picture, although with smaller statistics (cf. Table 2). We estimate that a ratio $S/\sqrt{S+B}=3$ (i.e. to reject the background-only hypothesis at $3\sigma$ level) could only be achieved in $10$ yr at this detector, if $f_{\text{MR}}\gtrsim 0.27$–assuming $100\%$ detection efficiency and the backgrounds from Ref. [83]. Note that these sensitivity prospects rely on optimistic detector characteristics (e.g. perfect detection efficiency and a detector configuration foreseeing four tanks of $10$ kt each, whereas only the deployment of two tanks is envisioned for DUNE Phase I [84]). The difference between the mean energies of electron and non-electron flavor neutrinos is larger than the one for antineutrinos. This could potentially disentangle some of the existing uncertainties in the picture, especially for what concerns the flavor conversion physics and the fraction of neutrino-driven BH-forming collapses [29]. However, even under optimistic assumptions, the contribution of DUNE to this matter is likely to be marginal.
@@ VII Conclusions and outlook
As the detection of the DSNB approaches, it is timely to address whether the diffuse neutrino emission from magnetorotational core collapses, leading to the formation of protomagnetars or spinars, could contribute to the DSNB. We rely on a suite of seven 3D special relativistic neutrino-magnetohydrodynamic simulations with different ZAMS mass, inspired by the 2D equivalent models presented in Ref. [40], to investigate the related neutrino properties and the associated contribution to the DSNB. Our models predict large neutrino luminosities $\gtrsim 10^{52}$ $\cdot$ s-1 for several seconds and large mean energies for non-electron neutrinos and antineutrinos ($\langle\varepsilon_{\nu_{x}}\rangle\gtrsim 20$ MeV). Such neutrino properties suggest that magnetorotational core collapses could provide a non-negligible contribution to the high-energy DSNB tail (above $5$ MeV). In this work, the evolution of the magnetorotational core collapses forming a protomagnetar is computed up to several seconds. A longer simulation would be desiderable to determine the evolution of the neutrino emission properties and assess the impact of the late-time emission on the diffuse emission and the potential numerical influence on the gradual rise of the mean energies at late times. The approach here adopted ensures that our results are conservative and do not overpredict the contribution of protomagnetars to the diffuse flux from MR core collapses.
@@
Given its larger size, the upcoming Hyper-Kamiokande-Gadolinium neutrino telescope could measure a fraction of magnetorotational collapses larger than $13\%$ at $3\sigma$ after $20$ yr of data-taking, if meanwhile we can infer the fraction of (neutrino-driven) BH-forming collapses and pin down the uncertainties on the CCSN rate through electromagnetic observations. This measurement would further benefit from improved background reduction and detection efficiency.